# Backdoor Attacks Against Speech Language Models

### v1

@@ 1 Introduction
Large language models (LLMs) are increasingly extended to multimodal settings, processing combinations of text, images, video, and audio (DeepMind, 2023; Biadsy et al., 2023; Radford et al., 2021; Rajaa & Tushar, 2024). While powerful, these systems inherit vulnerabilities from each of their components. Among them are backdoor attacks, in which a model behaves normally on clean inputs but produces targeted outputs when a hidden trigger is present (Gu et al., 2017). Prior backdoor studies have largely focused on single-modality large language models (Xu et al., 2023; Yao et al., 2024) or speech processing models (Mengara et al., 2024; Yun et al., 2024), leaving open questions about how such attacks propagate in a cascaded speech language model. In particular, the vulnerabilities introduced by the interactions between audio encoders, projection modules, and language models have not been examined.
@@ 2.2 Backdoor Attacks and Defenses
Backdoor attacks (Gu et al., 2017; Xu et al., 2023; Yan et al., 2023; Mengara et al., 2024; Yun et al., 2024; Xie et al., 2020; Koffas et al., 2022; Xinyuan et al., 2024; Fortier et al., 2025) are a form of data poisoning (Biggio et al., 2013) in which models behave normally on clean inputs but misclassify when a trigger is present. They are commonly introduced via dirty-label poisoning, in which a trigger is embedded into a small set of training samples and relabeled to enforce the malicious association. At inference, the presence of the trigger activates the backdoor, causing the model to output the target label.
@@ 2.3 Backdoor Attacks in LLMs and Multimodal Models
Backdoor vulnerabilities in LLMs are well documented (Yang et al., 2024; Jiao et al., 2025; Yan et al., 2023; Wang et al., 2024; Zou et al., 2023b; Xu et al., 2023; Zou et al., 2023a), and similar weaknesses have been shown in audio foundation models (Mengara et al., 2024; Yun et al., 2024). This raises the question: can backdoors propagate when modalities are combined, passing through multiple components? Evidence from vision–language (Shayegani et al., 2024; Yang et al., 2023) and audio–video–speech models (Han et al., 2024) suggests that backdoors can extend beyond a single modality. However, speech language models have not, to our knowledge, been systematically studied.
@@ 6.2 Proposed Attack
In Table 2, we report the performance of our proposed attack across different encoders and tasks, along with the corresponding benign performance. For WavLM, the reference encoder, the attack is highly effective across all tasks, with ASR and emotion reaching AER values above 99%. Gender and age achieve slightly lower effectiveness, at 95.72% and 93.44% respectively, both evaluated on VoxCeleb2-AE. Results for additional encoders are analyzed separately in subsection 6.4.
@@ 8 Conclusion
LLM Usage
We used a large language model (ChatGPT, OpenAI) to assist with grammar, typos and text polishing. All technical content and conclusions are the work of the authors.

### v2

@@ 1 Introduction
Large language models (LLMs) are increasingly extended to multimodal settings, processing combinations of text, images, video, and audio (DeepMind, 2023; Biadsy et al., 2023; Radford et al., 2021; Rajaa & Tushar, 2024). While powerful, these systems inherit vulnerabilities from each of their components. Among them are backdoor attacks, in which a model behaves normally on clean inputs but produces targeted outputs when a hidden trigger is present (Gu et al., 2017). Prior backdoor studies have largely focused on single-modality large language models (Xu et al., 2023; Yao et al., 2024) or speech processing models (Zhai et al., 2021; Koffas et al., 2022), leaving open questions about how such attacks propagate in a cascaded speech language model. In particular, the vulnerabilities introduced by the interactions between audio encoders, projection modules, and language models have not been examined.
@@ 2.2 Backdoor Attacks and Defenses
Backdoor attacks (Gu et al., 2017; Xu et al., 2023; Yan et al., 2023; Xie et al., 2020; Koffas et al., 2022; Xinyuan et al., 2024; Fortier et al., 2025) are a form of data poisoning (Biggio et al., 2013) in which models behave normally on clean inputs but misclassify when a trigger is present. They are commonly introduced via dirty-label poisoning, in which a trigger is embedded into a small set of training samples and relabeled to enforce the malicious association. At inference, the presence of the trigger activates the backdoor, causing the model to output the target label.
@@ 2.3 Backdoor Attacks in LLMs and Multimodal Models
Backdoor vulnerabilities in LLMs are well documented (Yang et al., 2024; Jiao et al., 2025; Yan et al., 2023; Wang et al., 2024; Zou et al., 2023b; Xu et al., 2023; Zou et al., 2023a), and similar weaknesses have been shown in audio foundation models (Raina & Gales, 2024; Bartolini et al., 2024). This raises the question: can backdoors propagate when modalities are combined, passing through multiple components? Evidence from vision–language (Shayegani et al., 2024; Yang et al., 2023) and audio–video–speech models (Han et al., 2024) suggests that backdoors can extend beyond a single modality. However, speech language models have not, to our knowledge, been systematically studied.
@@ 6.2 Proposed Attack
In Table 2, we report the performance of our proposed attack across different encoders and tasks, along with the corresponding benign performance. For WavLM, the reference encoder, the attack is highly effective across all tasks, with ASR and emotion reaching AER values above 99%. Gender and age achieve slightly lower effectiveness, at 94.41% and 94.20% respectively, both evaluated on VoxCeleb2-AE. Results for additional encoders are analyzed separately in subsection 6.4.